the altitude is 5 less than the base. the area is 42 sq.inches. find the base and altitude
Answer by mananth(16949)

let the base be x
altitude = x-5
Area of triangle = 1/2 * base * altitude
=1/2 *x*(x-5) = 42
x^2-5x=84
x^2-5x-84=0
x^2-12x+7x-84=0
x(x-12)+7(x-12)=0
(x-12)(x+7)=0
x=12 OR x=-7
So the base is 12
Altitude = 12-5=7